OpenStudy (anonymous):

Charles law- Physical Science The temperature inside my refrigerator is about 4 c. If I place a balloon in my fridge that initially has a temperature of 22 c and a volume of 0.5 liters, what will be the volume of the balloon when it is fully cooled by my refrigerator? I NEED HELP!

OpenStudy (deepolisnoob):

\[\frac{ V_{i} }{ T_{i} }=\frac{ V_{f} }{ T_{f} }\] You plug in your initial volume and temperature into the first part of the equation. You know your final temperature, too, which is 4 c. You must convert your C to K when working with the gas laws, so you add 274.15 to your C to turn it to a K. \[\frac{ 0.5l }{ 296.15K }=\frac{ V_{f} }{ 278.15K }\] You then try to get your V final by itself, so you multiply both sides by 278.15K, \[\frac{ 0.5l*278.15K }{ 296.15K }=V_{f}\] You K cancel in the numerator, so the only term you'll have left is l. When you simplify the numerator, you'll have . \[\frac{ 139.075l }{ 296.15 }=V_{f}\] Divide 139.075l by 296.15, and your new volume will be 0.47l. You can verify this with: \[\frac{ 0.5l }{ 296.15 }=\frac{ 0.47l }{ 278.15 }\] One mistake I just remembered is that you need to convert your Celcius to Kelvin, so I had to replace all C with K after I thought I finished :P

OpenStudy (deepolisnoob):

In the final equation, I forgot to add K to the end of the temperatures in the denominators